Please help me to solve this problem: 3 + lnX = 8
It would be greatly appreciated.
*Yasmin

Answer by narayaba(40)

3 + lnX = 8

lnX = 8-3
lnX = 5 equation 1
lnX means log of x to the base e
lne is log of e to base e and this equals 1
lne = 1
thus lnX = 5* lne
lnX = ln(e^5)
since the log on the right and log on the left are to the same base
X = e^5